OpenStudy (anonymous):

Set B has 12 elements in it and set C has 7 elements in it Answer the following: 1. What is the maximum number of elements that could be in the intersection of sets B and C? 2. What is the minimum number of elements that could be in the intersection of sets B and C?

OpenStudy (anonymous):

My answer is #1. maximum: 7 minimum: 5

OpenStudy (anonymous):

Yes to the first, no to the second... what if they have no elements in common? :P

Parth (parthkohli):

1. Consider this: {1,1,1,1,1,1,1,1,1,1,1,1} and {1,1,1,1,1,1,1} 2. Consider this: {1,1,1,1,1,1,1,1,1,1,1,1} and {2,2,2,2,2,2,2}

Parth (parthkohli):

What @PeterPan said. Remember that the intersection is the set of common elements.

OpenStudy (anonymous):

so the minimum could be zero

OpenStudy (anonymous):

That's right :D

Parth (parthkohli):

Yay!

OpenStudy (anonymous):

thanks guys could you help me with another question

Parth (parthkohli):

Sure :-)

OpenStudy (anonymous):

NO >:( LOL JUST KIDDING GO AHEAD

OpenStudy (anonymous):

A birthday cake is in the shape of a right circular cylinder and with a diameter of 30cm and a height of 12cm. A piece in the shape of a sectoral wedge with central angle 22.5 degrees is cut from the center of the cake. Draw the cake and the piece cut from it. Then using II = 3.14, find: 1. What is the volume of the piece cut from the birthday cake? 2. What is the surface area of the piece cut from the birthday cake? 3. After the piece has been taken out, what is the volume of the remaining part of the cake? 4. After the piece has been taken out what is the surface area of the remaining part of the cake?

OpenStudy (anonymous):

@PeterPan and @ParthKohli ru guys still there?

OpenStudy (anonymous):

So many numbers o.O

OpenStudy (anonymous):

First, we might wanna know what the volume of the entire cake is... can you find it? :)

OpenStudy (anonymous):

V=II*r^2*h V= 3.14*15^2*12 V= 3.14*225*12 V = 8478 cm3

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

Okay :D Now, we wanna know what part of the cake has been taken away in the slice :D So, the angle it makes is 22.5 degrees? the whole cake is 360 degrees... What's 22.5 / 360?

OpenStudy (anonymous):

.0625

OpenStudy (anonymous):

That's right :) So the volume of the slice is this times the volume of the cake :D

OpenStudy (anonymous):

.0625*8478=529.875

OpenStudy (anonymous):

And that's the volume of your cake :D

OpenStudy (anonymous):

the slice, at least :D

OpenStudy (anonymous):

whats next how do we find surface area

OpenStudy (anonymous):

Surface area is a bit tricky... where to begin... Area of a sector? :D That's just a fancy name for area of a slice of a circle...

OpenStudy (anonymous):

whats the formula

OpenStudy (anonymous):

can you draw out the cake pls

OpenStudy (anonymous):

formula for the area of a sector is area of the circle times (angle of sector/360)

OpenStudy (anonymous):

44.15625

OpenStudy (anonymous):

steps?

OpenStudy (anonymous):

22.5/360*3.14*15^2

OpenStudy (anonymous):

Seems reasonable :D Now, there's a sector this size on top and at the bottom, right? so double this area :D

OpenStudy (anonymous):

88.3125

OpenStudy (anonymous):

Now, there are two rectangles, the sides of the slice, with length equal to the height of the cake and width equal to the radius. What's the area of that rectangle?

OpenStudy (anonymous):

where does the rectangle come from?

OpenStudy (anonymous):

hello @PeterPan

OpenStudy (anonymous):

Because there when you slice the cake, it looks like this